What is the total energy of an electron when its momentum is 10 times its it is multiplied by the light velocity (rest mass of electron equals to 9×10^-31 kg)
1
Expert's answer
2019-05-07T09:56:30-0400

The total energy of an electron depends on its rest energy and kinetic energy:


"E=\sqrt{(m_0c^2)^2+(pc)^2}=\sqrt{(m_0c^2)^2+([10m_0c]c)^2},"

"E=\sqrt{(9\cdot10^{-31}(3\cdot10^8)^2)^2+([10\cdot9\cdot10^{-31}\cdot3\cdot10^8]3\cdot10^8)^2}="

"=8.14\cdot10^{-13}\text{ J}."